# WCCNet: Wavelet-context Cooperative Network for Efficient Multispectral Pedestrian Detection

### v2

@@ IV-D Ablation Study & Further Analysis
Mid-Cat[13] 16.76 6.14 20.93 23.54 14.04
DMAF[12] 17.51 8.38 20.64 23.29 14.75
Inter-MA[16] 17.13 7.24 21.54 24.80 14.20
@@
17.16 8.63 20.20 22.85 13.84
17.18 8.66 19.53 22.34 12.65
17.49 10.45 19.90 22.54 14.36
17.50 10.44 19.05 21.17 13.36
Analysis of Crossmodal Rearranging Fusion
Quantitative Comparisons with Different Fusion Schemes. Table VII shows ablation studies on comparison between our proposed CMRF and other various fusion schemes. The different fusion schemes are all implemented within WCCNet framework. The Early Concatenation (Early-Cat) and Late Concatenation (Late-Cat) fusion schemes only concatenate multimodality feature maps at the first and the last stage of the backbone along the channel dimension respectively, and get the lowest precision. The Middle Concatenation (Mid-Cat) fusion scheme concatenates RGB and infrared features along the channel dimension at every stage of the backbone, and thus obtain higher accuracy with a minor decline in $MR^{-2}$. These channel-dimension concatenation-based fusion schemes neglect the interaction between multimodal features which is essential for extracting complementary information and filtering out redundant information, hence unsurprisingly they get the lowest precision.
The middle fusion schemes designed for crossmodal feature interaction including DMAF, Inter-MA, and CMRF, are also evaluated. It is worth mentioning that the features for middle fusion are all extracted after CE layer. The DMAF[12] and Inter-MA[16] both obtain complementary information by subtracting feature maps of different modalities along the channel dimension to extract numerically differential features. Instead, the proposed CMRF extracts complementary information through mining semantic-aware features in spatially relevant local regions shared by different modalities.
Daytime
(a) RGB
(b) Infrared
(c) Dynamic weight map
Night
Figure 8: RGB images, infrared images and dynamic weight maps of CMRF are shown in the above three columns respectively. These images are grouped by different times including daytime and night. The highlighted region in red color is the regions of interest for CMRF.
Among the above fusion schemes, CMRF achieves the highest accuracy with competitive inference time, outperforming the second-best DMAF by $7.7\%$ in accuracy.
Qualitative Comparisons with Different Fusion Schemes. We further conduct qualitative evaluation on different fusion schemes. The detection results of WCCNet compared with different fusion schemes are shown in Fig. 7, along with their attention maps of detection head after fusion. It can be seen that, in both day and night, WCCNet has the least amount of falsely detected or missed pedestrians. Even in complex scenarios where pedestrians are occluded or of small scale, WCCNet still locates them accurately.
@@
Analysis of Different Components of CMRF. The influence of different components of CMRF is investigated in Table VII. The baseline is set to that only Semantic Rearranging Fusion(SRF) is applied in WCCNet, while achieving $20.20$ $MR^{-2}$. The accuracy is improved to $19.53$ $MR^{-2}$ when Cross-domain Embedding(CE) layers are added to narrow the semantic gap of multimodal features before being fused by SRF. Lastly, the Cross-domain Spatial Alignment(CSA) is adopted, and then the accuracy further reaches $19.05$ $MR^{-2}$ as features of different spectra are aligned. From this, we can conclude that each component of CMRF indeed contributes to the final accuracy.
@@
Visual Illustration for CMRF. The effectiveness of CMRF is shown in Fig. 8, where the focused regions of context-specific features are visualized. The dynamic weight map in Fig. 8 is implemented by mapping local rearranging weight $\mathbf{W}_{l}$ of all localizations $l$ onto the corresponding images to get heatmaps. As shown in Fig. 8, even when the pedestrians are not clearly visible in any single modality, CMRF effectively focuses on the regions relevant to pedestrians. This observation demonstrates that the complementary features of pedestrians in spatially-related local regions for different modalities are effectively aggregated by CMRF.
V Conclusion
@@
In this paper, we have proposed a novel framework WCCNet to achieve efficient multispectral pedestrian detection for autonomous driving. Based on integrated wavelet with lower computational complexity, WCCNet has boosted the inference speed by learning multi-domain representations for different spectra while maintaining competitive accuracy. A novel fusion scheme CMRF has been proposed to merge the spatially relevant local regions shared by different spectra with adaptive weights learned on-the-fly. Extensive experiments on KAIST and FLIR datasets have been conducted and substantiate the aforementioned superiority of WCCNet over several state-of-the-art methods.

### v3

@@ 4.5.3 Analysis of Crossmodal Rearranging Fusion
Mid-Cat[MLPD2021] 16.76 6.14 20.93 23.54 14.04
DMAF[MBNet2020] 17.51 8.38 20.64 23.29 14.75
Inter-MA[zhang2023TripleNet] 17.13 7.24 21.54 24.80 14.20
@@
17.16 8.82 19.35 22.50 13.60
17.18 8.91 18.25 21.73 11.90
17.49 10.61 18.85 22.05 13.01
17.50 10.62 17.81 20.36 12.37
Table 6: Evaluation of different fusion settings for WCCNet, along with ablation study on the proposed components of CMRF. Accuracy is evaluated by $\mathrm{MR}^{-2}$ with All setting.
Analysis of Different Components in CMRF
@@
The influence of different components of CMRF is investigated in Table 6. The baseline is set to that only Semantic Rearranging Fusion(SRF) is applied in WCCNet, while achieving $20.20$ $MR^{-2}$. The accuracy is improved to $19.53$ $MR^{-2}$ when Cross-domain Embedding(CE) layers are added to narrow the semantic gap of multimodal features before being fused by SRF. Lastly, the Cross-domain Spatial Alignment(CSA) is adopted, and then the accuracy further reaches $19.05$ $MR^{-2}$ as features of different spectra are aligned. Each component of CMRF indeed contributes to the final accuracy.
Visual Illustration for CMRF
@@
As shown in Fig. 12, for difficult scenarios that pedestrians not clearly visible in any single modality or unrelated target interference, CMRF effectively focuses on the regions relevant to pedestrians. This observation demonstrates that the complementary features of pedestrians in spatially-related local regions for different modalities are effectively aggregated by CMRF.
5 Conclusion
@@
In this paper, we have proposed a novel framework WCCNet to achieve efficient multispectral pedestrian detection. Based on wavelet-cooperative backbone integrated with efficient MoWE layers, WCCNet has boosted the inference speed by learning spectra-specific representations while maintaining competitive accuracy. Based on environment context, MoWE leverages shared static DWT and selectively activated ADWT layers to efficiently extract rich and informative features from infrared images with low computational overhead. For further exploration on multispectral reciprocal cooperation, CMRF has been proposed to merge the spatially relevant local regions shared by different spectra with adaptive weights learned on-the-fly. Extensive experiments on KAIST and FLIR datasets have been conducted and substantiate the aforementioned superiority of WCCNet over several state-of-the-art methods. We believe that WCCNet can potentially benefit the development of resource-constrained detection systems for autonomous driving applications. In the future, we aim to further enhance the representation capability of MoWE and the fusion efficiency of CMRF, while extending the WCCNet framework to diverse backbones and broader multimodal tasks.